OpenStudy (anonymous):

The Sweet Candy Shop buys 600 pounds of chocolate covered strawberries at $5.59 per pound. If a 10% spoilage rate is anticipated, at what price per pound should the strawberries be sold in order to achieve a 60% markup based on cost?

OpenStudy (tkhunny):

Really, just think about it one piece at a time. No need to have the whole picture in your mind at once. Base Cost: 600*5.59 = 3,354.00 60% Markup on Base Cost: 0.60 * 3,354.00 = 2,012.40 Sales Goal: 3,354.00 + 2,012.40 = 5,366.40 Actual Sales: 600 - 0.10*600 = 600 - 60 = 540 Okay, how do you get $5,266.40 out of 540 lbs of strawberries.